Find the magnetic field at point P for the steady current configuration below.
Answer #1

The equation for the smaller coil is into the page

B_1= \frac{1}{4} \left ( \frac{\mu_o I}{2 a} \right )

B_1= \left ( \frac{\mu_o I}{8 a} \right )

The equation for the larger coil is

B_2= \frac{1}{4} \left ( \frac{\mu_o I}{2 b} \right )

B_2= \left ( \frac{\mu_o I}{8 b} \right )

net field is
B = B_1 - B_2=\left ( \frac{\mu_o I}{8 a} \right )- \left ( \frac{\mu_o I}{8 b} \right )

B = \left ( \frac{\mu_o I}{8 } \right ) \left ( \frac{1}{a} -\frac{1}{b} \right )

B = \left ( \frac{\mu_o I}{8 } \right ) \left ( \frac{b-a}{ab} \right )